OpenStudy (anonymous):

I need help with my math homework. It is over Equivelent Fractions. And I was wondering if you would help me with the problems:

OpenStudy (anonymous):

What are equivalent fractions?

OpenStudy (anonymous):

They are somthing that you use and they are the same fraction. Example. 6/12 and 12/24 is equvelent

OpenStudy (anonymous):

Why are they equivalent?

OpenStudy (anonymous):

They are equal they can go into each other

OpenStudy (anonymous):

Alright :-D

OpenStudy (anonymous):

I am guessing you cant help me. thank you for your time

OpenStudy (anonymous):

That's not really why they're equivalent.... 2 goes into 4.

OpenStudy (anonymous):

MetalPen can you help me with my homework then

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

The instructions say write a fraction equivelent to the given number. And the problem is 8. I dont understand that

OpenStudy (anonymous):

Give a fraction equivalent to the number 8? 8 is really 8/1.

OpenStudy (anonymous):

so it would be 8/1?

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

You understand why right?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Lol. Think of it just as 8 divided by 1, not as a fraction. What's 8 divided by 1? 8 right?

OpenStudy (anonymous):

Oh. I get it now. Thank you(: Do you mind helping me with an actual fraction?

OpenStudy (anonymous):

Sure. Oh and, since I think it's clear you are at a lower level, we could stay in contact and I could help you with your math since we will increase our knowledge in math monotonically.

OpenStudy (anonymous):

okayy... the problem is 2 2/3

OpenStudy (anonymous):

Okay. Well turn it into an improper fraction first.

OpenStudy (anonymous):

2=2/1, and 2/1 + 1/3 is what?

OpenStudy (phi):

8= 16/2 = 24/3 and of course lots of other examples

OpenStudy (anonymous):

6+2=8 so 8/3

OpenStudy (anonymous):

Yup.

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Lol. Anymore questions?

OpenStudy (anonymous):

So all you have to do is come up with your own fractions that 'go into' the given number.

OpenStudy (anonymous):

yess, just to make sure I am right

OpenStudy (anonymous):

9 8/11 so 9*11= 99+8=107 so it is 107/11

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

Yup.

OpenStudy (anonymous):

Okay. Thank you(: i gave like 3 more lol

OpenStudy (anonymous):

*i got 3 more

OpenStudy (anonymous):

Okay sure.

OpenStudy (anonymous):

4 26/13

OpenStudy (anonymous):

78/13

OpenStudy (anonymous):

Am I correct? I did it in my head.

OpenStudy (anonymous):

So it might not be accurate. :(

OpenStudy (anonymous):

I got 68/13

OpenStudy (anonymous):

is the fraction \( 4 \frac{26}{13}\) ?

OpenStudy (anonymous):

Let's see. Rounded it's about 40+30.

OpenStudy (anonymous):

or is it \( \frac{6}{13} \)

OpenStudy (anonymous):

So I'll have to do the math out.

OpenStudy (anonymous):

yes, first one ag

OpenStudy (anonymous):